# Jizhou Liu Booth School of Business

### v1

@@ 4 Randomization Tests for Weak Null Hypotheses
    Boundedness: for all buyer-seller pairs $(i,j)$ and all $b,s\in\{0,1\}$, $|Y_{i,j}(b,s)|\leq C_{2}$.
For any level $\alpha\in(0,1)$, the testing procedure in Procedure 4.1 satisfies:
under the null hypothesis $H_{0}^{wb}$, where the expectation is with respect to $p(\mathbf{W}\mid\mathcal{C})$.
The proof of Theorem 4.1 utilizes a common strategy in the literature, demonstrating that under the studentized test statistic the sampling distribution is stochastically dominated by the randomization distribution. The analysis begins by considering the case conditional on a fixed seller-side randomization $w^{S}$, then reducing the problem to a single-population scenario. In this setup, unit-level potential outcomes for a given buyer are defined as the average over untreated sellers, with randomization solely applied to the buyer side. Leveraging existing results (Zhao and Ding, 2021; Wu and Ding, 2021), the studentized test statistic yields a randomization distribution that converges asymptotically to a standard normal distribution that stochastically dominates the sampling distribution. A primary challenge lies in extending this result to the unconditional scenario as $I,J\rightarrow\infty$. A crucial step is to observe that the conditional sampling distribution is centered at the point
$\tau(w^{S})=\frac{1}{IJ_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(1,0)(1-w^{S}_{j})-\frac{1}{IJ_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(0,0)(1-w^{S}_{j}),$
whereas the weak null hypothesis concerns $\tau$, given by
$\tau=\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(0,0)-\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(1,0).$
Following the methodological approach of Masoero et al. (2024), we demonstrate the concentration of the conditional mean and variance, thereby substantiating the validity.
Remark 4.1.
The procedure for assessing total effects should adhere to that described in Procedure 4.1, substituting $\bar{Y}_{i}^{B}$ with the block average. Section 3.3.2 advises selecting the maximal permissible block size $k$, constrained by a predetermined maximum power level. However, for the weak null hypothesis, a large number of blocks is essential to ensure the accuracy of the variance estimator $V^{B}(\mathbf{W}\mid\mathbf{Y},\mathcal{C})$. The simulation results from Table 4 illustrate deteriorating size control with increasing block size, necessitating a smaller block size $k$ for testing the weak null hypothesis.
5 Simulations
In this section, we examine the finite-sample behavior of our randomization tests for both total and spillover effects. Particularly, we compare the performance of randomization test against Neymanian style inference proposed in Masoero et al. (2024). Suppose that there are $3n$ units in each population, i.e. $I=J=3n$, and the design is a completely randomized two-sided design with $n$ units assigned to treatment group in both populations. Following Section 7.1 of Masoero et al. (2024), potential outcomes are generated as follows.
$Y_{ij}(w,h)\stackrel{{\scriptstyle\text{ ind }}}{{\sim}}\begin{cases}F_{0}(\cdot)&\text{ if }w=h=0,\\ F_{0}(\cdot)+F_{B}(\cdot)&\text{ if }w=1,h=0,\\ F_{0}(\cdot)+F_{S}(\cdot)&\text{ if }w=0,h=1,\\ F_{0}(\cdot)+F_{1}(\cdot)&\text{ if }w=h=1,\end{cases}$
where $F_{\ell}$ are normal distributions such that $F_{\ell}(\cdot)=\mathcal{N}\left(\mu_{\ell},\sigma_{\ell}^{2}\right)$ for $\ell\in\{0,B,S,1\}$. In such a setting, we consider the problem of testing row spillover and total treatment effects, and compare our testing procedures from Procedure 3.1, 3.2 and 4.1 with Masoero et al. (2024). Under the sharp null hypothesis, parameters are set as follows: $\mu_{0}=\mu_{S}=\mu_{B}=\mu_{1}=0$, $\sigma_{0}=0.2$ and $\sigma_{B}=\sigma_{S}=\sigma_{1}=0$. Under the weak null hypothesis, $\mu_{l}$ are set to be the same as under sharp null, but variances are set to be different: $\sigma_{0}=0.2,\sigma_{B}=\sigma_{1}=0.4$ and $\sigma_{S}=0$. Under the alternative hypotheses, we set the parameters in a similar way but set the spillover effect to be 0.01 and total effects to be 0.02, i.e. $\mu_{B}=0.01,\mu_{1}=0.02$.
Under $H_{0}$ Under $H_{1}$ Procedure $n$ FRT Neymanian FRT adjusted FRT Neymanian FRT adjusted 10 5.02% 0.68% 4.78% 8.88% 3.84% 8.84% Test Buyer 20 4.82% 3.04% 5.26% 19.96% 11.92% 20.46% Spillover Effects 30 5.00% 3.66% 4.90% 42.42% 28.24% 41.94% (Procedure 3.1) 40 4.78% 2.86% 5.02% 58.50% 52.40% 57.76% 50 4.72% 2.42% 4.62% 81.42% 73.66% 81.26% 100 4.74% 3.28% 4.84% 100.00% 100.00% 100.00% 10 4.78% 1.82% 5.02% 5.64% 3.66% 5.50% Test 20 4.84% 1.24% 4.94% 10.16% 23.32% 10.26% Total Effects 30 4.84% 0.92% 4.70% 18.92% 57.52% 17.44% (Procedure 3.2) 40 4.96% 1.60% 5.12% 30.58% 86.56% 30.04% 50 5.06% 1.56% 4.82% 42.58% 97.90% 40.34% 100 5.24% 1.28% 4.92% 95.38% 100.00% 93.82%
Table 1 displays the rejection probabilities under the sharp null and alternative hypotheses, computed from 5,000 Monte Carlo replications with the $p$-values approximated by 500 independent permutations of the treatment vector in each replication. “FRT” stands for the testing procedure based on sharp null hypotheses as in Procedure 3.1 and 3.2. “FRT adjusted” stands for the studentized randomization tests as in Procedure 4.1. “Neymanian” stands for $t$-tests using conservative estimators of variances from Masoero et al. (2024). The block size for Procedure 3.2 is set to be $k=\lfloor n/4\rfloor$, resulting in $|\mathcal{W}|=C(12,4)=495$ and a maximum power approximately equal to 0.955. The results show that the rejection probabilities of both “FRT” and “FRT adjusted” are universally around 0.05 under the null hypothesis, which verifies the finite-sample exactness of our tests across all the designs. Meanwhile, the Neymanian inference method is conservative as expected by Masoero et al. (2024). Under the alternative hypotheses, the rejection probabilities of our methods are higher than Neymanian method when testing spillover effects, but lower when testing total effects. The power gain by Procedure 3.1 likely comes from the test’s exactness, whereas the power loss from Procedure 3.2 is due to the loss in sample size used for calculating test statistics.
Block Size $k$ Hypothesis Method 1 2 4 5 10 20 25 50 $H_{0}^{total}$ FRT 5.22% 5.40% 5.34% 5.04% 4.90% 4.98% 4.88% 0.84% FRT adjusted 5.24% 4.84% 4.88% 4.86% 5.26% 4.98% 5.02% 1.08% $H_{1}^{total}$ FRT 13.20% 19.56% 35.90% 43.46% 69.40% 91.06% 95.10% 6.98% FRT adjusted 13.32% 20.78% 35.92% 41.80% 69.28% 89.68% 93.56% 13.62%
We further investigate the robustness of the randomization inference by examining the influence of block size on both the size and power of the testing procedures. Table 2 presents our analysis of rejection probabilities under the sharp null and alternative hypotheses, following the same Monte Carlo setup in Table 1. We observe that the statistical power of the methods under investigation increases monotonically with the enlargement of the block size. Specifically, when block size is set at $k=25$, where the randomization space encompasses $|\mathcal{W}|=C(12,4)=495$, near-maximum power is achieved. However, a further increase in block size to $k=50$ results in a diminished randomization space of $|\mathcal{W}|=C(6,2)=15$, leading to a notable decline in test power. These results are consistent with the pattern observed in the graphical illustration of the theoretical lower bound in Figure 3.
5.1 Weak null hypotheses
Table 3 presents the rejection probabilities under the weak null and alternative hypotheses, also computed from 5,000 Monte Carlo replications and 500 independent permutations of the treatment vector in each replication. In Table 3, the block size is fixed to be $2$, so that the number of block is large enough to guarantee the precision of variance estimation.
Under $H_{0}$ Under $H_{1}$ Procedure $n$ FRT Neymanian FRT adjusted FRT Neymanian FRT adjusted 10 9.12% 0.6% 4.24% 14.08% 3.34% 8.20% Test Buyer 20 9.60% 2.24% 4.30% 16.20% 1.94% 8.56% Spillover Effects 30 9.54% 1.26% 3.98% 22.84% 8.46% 12.86% (Procedure 4.1) 40 8.50% 1.80% 3.36% 35.90% 15.00% 21.06% 50 9.62% 2.06% 4.20% 46.74% 23.14% 31.24% 100 8.64% 1.26% 3.00% 93.86% 82.98% 87.70% 10 13.84% 0.82% 8.22% 15.24% 1.36% 8.16% Test 20 12.14% 1.18% 6.04% 16.48% 5.36% 10.64% Total Effects 30 11.90% 1.30% 5.70% 19.78% 10.48% 13.36% (Procedure 4.1) 40 11.46% 0.84% 5.16% 24.76% 22.16% 16.74% 50 11.56% 1.18% 5.52% 30.74% 38.04% 21.30% 100 11.44% 0.86% 4.84% 67.32% 97.82% 53.56%
Table 3: Rejection probabilities under weak null and alternative hypothesis
In Table 4, we characterize the effect of block size on the performance of total effect tests by following the same simulation study in Table 2. Our findings here illuminate a distinctive pattern: under the weak null hypothesis, the standard Fisher Randomization Test (FRT) tends to over-reject, whereas the adjusted FRT remains valid. Notably, the spillover effect tests are conservative, while the total effect tests are close to exactness when the block size is small and invalid when the block size is too large. The asymptotic exactness of total effect tests is potentially due to the effect of “sampling-based versus design based uncertainty” (Abadie et al. (2020)). The data generating process can be thought of as subsampling the focal units $|\mathcal{U}|$ from the finite population of the entire matrix. In the spillover effect tests, $|\mathcal{U}|=O(n^{2})$, while the total effect tests have $|\mathcal{U}|=O(n)$. Therefore, the spillover effect tests are more subject to the design-based finite population framework, while the total effect tests are closer to the sampling-based super population framework.
Block Size $k$ Hypothesis Method 1 2 4 5 10 20 25 50 $H_{0}^{total}$ FRT 11.22% 12.24% 11.68% 12.10% 13.42% 12.76% 13.24% 1.52% FRT adjusted 5.00% 5.32% 5.04% 5.82% 5.74% 8.12% 8.44% 2.04% $H_{1}^{total}$ FRT 14.94% 18.48% 24.58% 27.40% 40.62% 60.34% 66.48% 6.78% FRT adjusted 7.32% 9.60% 14.36% 16.90% 27.44% 44.46% 54.44% 9.34%
Table 4: Rejection probabilities under different block sizes with $n=100$ under weak null
@@ 6 Conclusion
Appendix A Futher Details
@@
Let $n=|\mathcal{U}|$ and $m=|\mathcal{Z}|$. Let the randomization distribution and the null distribution be denoted, respectively, by
$T(\mathbf{W}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})\sim\hat{F}_{1,n,m},\text{ and }T(\mathbf{W}\mid\mathbf{Y}^{obs},\mathcal{C})\sim\hat{F}_{0,n,m},\text{ where }\mathbf{W}\sim P(\mathbf{W}\mid\mathcal{C}).$ (9)
@@ B.1.1 Proof for Randomization Test
The proof is a direct application of Theorem 1 of Basse et al. (2019) by verifying that any test statistic restricted on $\mathcal{C}$ is imputable, i.e. $T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C})=T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})$. For clarity in exposition, we first show imputability for the difference-in-mean estimators in Procedure 3.1 and 3.2, and then repeat the same argument for a general test statistics. Note that, in Procedure 3.1
$T(\mathbf{W}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})=\frac{1}{n_{1}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},w^{S}_{j})(1-w^{S}_{j})w^{B}_{i}-\frac{1}{n_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},w^{S}_{j})(1-w^{S}_{j})(1-w^{B}_{i})~{}.$
Consider $\mathbf{W}^{\prime}=w^{B,\prime}(w^{S,\prime})^{\top}\in\mathcal{W}$. Under the conditioning set of assignments $\mathcal{W}=\{v(w^{S})^{\top}:v\in\mathbb{W}^{B}\}\cap\mathcal{M}$, we have $w^{S,\prime}=w^{S}$. Then,
@@
$=\frac{1}{n_{1}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},w_{j}^{S})(1-w^{S}_{j})w^{B,\prime}_{i}-\frac{1}{n_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},w_{j}^{S})(1-w^{S}_{j})(1-w^{B,\prime}_{i})$
$=\frac{1}{n_{1}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},0)(1-w^{S}_{j})w^{B,\prime}_{i}-\frac{1}{n_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},0)(1-w^{S}_{j})(1-w^{B,\prime}_{i})~{}.$
Under $H_{0}^{buyer}$, we have
$Y_{i,j}(1,0)=Y_{i,j}(0,0)\text{ for all }i\in[I],j\in[J]~{}.$
Therefore,
$T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})=\frac{1}{n_{1}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w^{B,\prime}_{i},0)(1-w^{S}_{j})w^{B,\prime}-\frac{1}{n_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w^{B,\prime},0)(1-w^{S}_{j})(1-w^{B,\prime}_{i})$
$=T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C})~{}.$
Next, we show imputability for Procedure 3.2. Note that
$T(\mathbf{W}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C}^{(k)})=\frac{1}{n_{1,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{S}_{j})w^{S}_{j}w^{B}_{i}-\frac{1}{n_{0,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{S}_{j})(1-w^{S}_{j})(1-w^{B}_{i})~{}.$
Under the conditioning event, we have $w_{i}^{B}=w_{j}^{S}$ for all $(i,j)\in\mathcal{U}$. Therefore, we have
$T(\mathbf{W}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C}^{(k)})=\frac{1}{n_{1,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{B}_{i})w^{B}_{i}-\frac{1}{n_{0,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{B}_{i})(1-w^{B}_{i})~{}.$
Then, by $H_{0}^{total}:Y_{i,j}(1,1)=Y_{i,j}(0,0)\text{ for all }i\in[I],j\in[J]$,
$T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C}^{(k)})$ $=\frac{1}{n_{1,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{B}_{i})w^{B,\prime}_{i}-\frac{1}{n_{0,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{B}_{i})(1-w^{B,\prime}_{i})$
$=\frac{1}{n_{1,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B,\prime}_{i},w^{B,\prime}_{i})w^{B,\prime}_{i}-\frac{1}{n_{0,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B,\prime}_{i},w^{B,\prime}_{i})(1-w^{B,\prime}_{i})$
$=T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C}^{(k)})~{}.$
Now, we show imputability for general test statstics in Procedure 3.1. Note that
$T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})$ $=T(\mathbf{W}^{\prime}\mid\{\mathbf{Y}_{i,j}(w_{i}^{B},w_{j}^{S})\}_{(i,j):w_{j}^{S}=0},\{w^{S,\prime}=w^{S}\},\mathcal{C})$
$=T(\mathbf{W}^{\prime}\mid\{\mathbf{Y}_{i,j}(w_{i}^{B},0)\}_{(i,j):w_{j}^{S}=0},\{w^{S,\prime}=0\},\mathcal{C})$
$=T(\mathbf{W}^{\prime}\mid\{\mathbf{Y}_{i,j}(w_{i}^{B,\prime},0)\}_{(i,j):w_{j}^{S,\prime}=0},\{w^{S,\prime}=0\},\mathcal{C})\quad\quad\text{(by $H_{0}^{total}$)}$
$=T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C})~{}.$
In Procedure 3.2,
$T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})$ $=T_{k}(\mathbf{W}^{\prime}\mid\{\mathbf{Y}_{i,j}(w_{i}^{B},w_{j}^{S})\}_{(i,j)\in\mathcal{U}},\{w_{i}^{B,\prime}=w_{j}^{S,\prime}\},\mathcal{C}^{(k)})$
$=T_{k}(\mathbf{W}^{\prime}\mid\{\mathbf{Y}_{i,j}(w_{i}^{B},w_{i}^{B})\}_{(i,j)\in\mathcal{U}},\{w_{i}^{B,\prime}=w_{j}^{S,\prime}\},\mathcal{C}^{(k)})\quad\quad\quad\text{($w_{i}^{B}=w_{j}^{S}$ for $(i,j)\in\mathcal{U}$)}$
$=T_{k}(\mathbf{W}^{\prime}\mid\{\mathbf{Y}_{i,j}(w_{i}^{B,\prime},w_{i}^{B,\prime})\}_{(i,j)\in\mathcal{U}},\{w_{i}^{B,\prime}=w_{j}^{S,\prime}\},\mathcal{C}^{(k)})\quad\quad\quad\text{(by $H_{0}^{total}$)}$
$=T^{B}(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C}^{(k)})~{}.$

### v2

@@ 6 Data Application
    $H_{0}^{(0,1)}:Y_{i,j}(0,0)=Y_{i,j}(0,1)$ for all $i,j$ with $X_{i}=1$ and $X_{j}=0$, which tests whether treating the less constrained household (seller) alters its connection with an untreated, constrained household (buyer).
  3. 3.
    $H_{0}^{(1,1)}:Y_{i,j}(0,0)=Y_{i,j}(1,1)$ for all $i,j$ with $X_{i}=1$ and $X_{j}=0$, which tests whether jointly treating both households changes their financial connection relative to the case where neither is treated.
Although the original experiment was not designed as a two-sided market intervention, the setting naturally fits our framework. The outcome of interest is dyadic—measured at the household-pair level—so the experimental design can be interpreted as a two-sided randomization problem where the “buyer” and “seller” sides originate from the same population. In this setting, the outcome $\mathbf{Y}$ is a $915\times 915$ matrix and the assignment vector is 915‑dimensional, with the buyer‑side and seller‑side assignments coinciding, i.e., $w^{B}=w^{S}$. In our analysis, we take a further step to partition households into “buyers” and “sellers” using baseline information on which households experienced shocks. This not only allows us to more directly apply our two-sided randomization framework but also reflects a meaningful economic distinction that enables analysis of treatment effect heterogeneity.
To test $H_{0}^{(1,0)}$ and $H_{0}^{(0,1)}$, we apply the one-sided permutation procedure described in Procedure 3.1. For $H_{0}^{\text{total}}$, we leverage the fact that the network is censored as survey data record only within-village links. This allows us to use the block-wise permutation procedure in Procedure 3.2, where each village defines a permutation block. Table 4 reports $p$-values from our randomization-based tests alongside those from a standard two-sample $t$-test. [^7] We omit the results for “FRT two-way” because they are more conservative than those for FRT adjusted” and are therefore not statistically significant. We include the $t$-test for comparison, noting that the Neymanian-style methods (Masoero et al., 2026) are not directly applicable due to the censored nature of the data. In contrast, our approach remains valid by conditioning on the observed household pairs.
Hypothesis FRT FRT adjusted Two-sample $t$-test $Y_{i,j}(0,0)=Y_{i,j}(1,0)$ 0.49 0.50 0.49 $Y_{i,j}(0,0)=Y_{i,j}(0,1)$ 0.42 0.41 0.41 $Y_{i,j}(0,0)=Y_{i,j}(1,1)$ 0.70 0.69 0.77
Table 4: $p$-values on testing the treatment effects on financial network links.
We fail to reject all three null hypotheses, indicating that providing savings accounts did not significantly affect households’ risk-sharing relationships. This finding complements Comola and Prina (2021), who document significant effects on network behavior in the full population, which suggests that such effects are not driven by households experiencing shocks. Consistent with prior evidence, the impacts of savings accounts appear highly context-dependent: some studies find limited usage among poor households (Dupas et al., 2016), while others report high take-up but no clear effects on aggregate expenditure, assets, or income (Prina, 2015). Overall, these results suggest that the effectiveness of savings accounts depends critically on how households use them, and that access alone may be insufficient to change financial behavior.
@@ 7 Conclusion
Appendix A Further Details
@@
Let $n=|\mathcal{U}|$ and $m=|\mathcal{W}|$. Let the randomization distribution and the null distribution be denoted, respectively, by
[…]
@@ B.1.1 Proof for Randomization Test
The proof is a direct application of Theorem 2 of Puelz et al. (2021). We first verify that, within the conditioning event $\mathcal{C}$, the test statistic is imputable under the null: for any $\mathbf{W}^{\prime}\in\mathcal{W}(\mathcal{C})$,
$T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C})=T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C}).$
We first show imputability for the difference-in-means statistics in Procedure 3.1 and Procedure 3.2, and then for a generic statistic.
Procedure 3.1.
Recall
$T(\mathbf{W}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})=$ $\ \frac{1}{n_{1}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},w^{S}_{j})\,(1-w^{S}_{j})\,w^{B}_{i}$
$-\frac{1}{n_{0}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},w^{S}_{j})\,(1-w^{S}_{j})\,(1-w^{B}_{i}).$
Fix any $\mathbf{W}^{\prime}=w^{B,\prime}(w^{S,\prime})^{\top}\in\mathcal{W}(\mathcal{C})$. Under the conditioning event for Procedure 3.1, the seller-side assignment is fixed, so $w^{S,\prime}=w^{S}$. Hence,
@@
$=$ $\ \frac{1}{n_{1}}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{i,j}(w_{i}^{B},w^{S}_{j})\,(1-w^{S}_{j})\,w^{B,\prime}_{i}$
[…]
where the last equality uses that $(1-w_{j}^{S})$ restricts attention to pairs with $w_{j}^{S}=0$.
Under $H_{0}^{\text{buyer}}$, we have $Y_{i,j}(1,0)=Y_{i,j}(0,0)$ for all $i,j$, so for any $z\in\{0,1\}$, $Y_{i,j}(z,0)=Y_{i,j}(0,0)$. Therefore we may replace $Y_{i,j}(w_{i}^{B},0)$ by $Y_{i,j}(w_{i}^{B,\prime},0)$:
$T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C})$
[…]
$=$ $\ T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C}),$
which establishes imputability.
Procedure 3.2.
Recall
$T(\mathbf{W}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C}^{(k)})=\frac{1}{n_{1,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{S}_{j})\,w^{S}_{j}w^{B}_{i}-\frac{1}{n_{0,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{S}_{j})\,(1-w^{S}_{j})(1-w^{B}_{i}).$
Under $\mathcal{C}^{(k)}$, we have $w_{i}^{B}=w_{j}^{S}$ for all $(i,j)\in\mathcal{U}$, so
$T(\mathbf{W}\mid\mathbf{Y}(\mathbf{W}),\mathcal{C}^{(k)})=\frac{1}{n_{1,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{B}_{i})\,w^{B}_{i}-\frac{1}{n_{0,k}}\sum_{(i,j)\in\mathcal{U}}Y_{i,j}(w^{B}_{i},w^{B}_{i})\,(1-w^{B}_{i}).$
Fix $\mathbf{W}^{\prime}\in\mathcal{W}(\mathcal{C}^{(k)})$. Then $\mathbf{W}^{\prime}$ also satisfies the same block constraint, hence for all $(i,j)\in\mathcal{U}$, $w_{i}^{B,\prime}=w_{j}^{S,\prime}$ and in particular we may write $Y_{i,j}(w_{i}^{B,\prime},w_{i}^{B,\prime})$. By $H_{0}^{\text{total}}$, $Y_{i,j}(1,1)=Y_{i,j}(0,0)$ for all $i,j$, so $Y_{i,j}(z,z)$ is constant across $z\in\{0,1\}$. Thus we can replace $Y_{i,j}(w_{i}^{B},w_{i}^{B})$ by $Y_{i,j}(w_{i}^{B,\prime},w_{i}^{B,\prime})$ inside $T(\mathbf{W}^{\prime}|\cdot)$:
[…]
$=$ $\ T(\mathbf{W}^{\prime}\mid\mathbf{Y}(\mathbf{W}^{\prime}),\mathcal{C}^{(k)}),$
establishing imputability.
Generic statistics.
The same reasoning applies to any statistic that depends on $(\mathbf{Y}(\mathbf{W}^{\prime}),\mathbf{W}^{\prime})$ only through outcomes of pairs whose exposure is fixed under $\mathcal{C}$.
For Procedure 3.1, $\mathcal{C}$ fixes $w^{S}$, hence fixes the set $\{(i,j):w_{j}^{S}=0\}$. Under $H_{0}^{\text{buyer}}$, all potential outcomes $Y_{i,j}(z,0)$ coincide across $z\in\{0,1\}$, so for any $\mathbf{W}^{\prime}\in\mathcal{W}(\mathcal{C})$,
$\{Y_{i,j}(w_{i}^{B},0)\}_{(i,j):w_{j}^{S}=0}=\{Y_{i,j}(w_{i}^{B,\prime},0)\}_{(i,j):w_{j}^{S,\prime}=0}.$
Therefore any such statistic is imputable (note: the null used here is $H_{0}^{\text{buyer}}$, not $H_{0}^{\text{total}}$).
For Procedure 3.2, $\mathcal{C}^{(k)}$ fixes the constraint $w_{i}^{B}=w_{j}^{S}$ for $(i,j)\in\mathcal{U}$ and under $H_{0}^{\text{total}}$ the values $Y_{i,j}(z,z)$ coincide across $z$, so any statistic depending on $\{Y_{i,j}(w_{i}^{B},w_{j}^{S})\}_{(i,j)\in\mathcal{U}}$ through $\{Y_{i,j}(w_{i}^{B},w_{i}^{B})\}_{(i,j)\in\mathcal{U}}$ is imputable by the same substitution.
Correct randomization distribution.
We show that Procedures 3.1 and 3.2—with Procedure 3.2 replaced by the conditional distribution described in Remark 3.2 and (7)—induce randomization distributions consistent with
$p(\mathbf{W}\mid\mathcal{C})\;\propto\;p(\mathcal{C}\mid\mathbf{W})\,p(\mathbf{W}).$
Following the argument in the proof of Theorem 2 of Puelz et al. (2021), it suffices to show that, for every $\mathbf{w}\in\mathcal{W}$,
$p(\mathcal{C}\mid\mathbf{W}=\mathbf{w})=\mathbb{I}\{\mathbf{w}\in\mathcal{W}\}.$
This condition holds for Procedure 3.1 by construction.
For Procedure 3.2, the conditioning mechanism does not favor one focal assignment over another. Indeed, the partitions $\mathcal{I}^{(k)}(\mathbf{w})$ and $\mathcal{J}^{(k)}(\mathbf{w})$ are themselves generated at random; equivalently, one may independently shuffle the buyer and seller indices within the treated and control groups and then form blocks of size $k$ sequentially. This construction is uniform over all assignments compatible with the partition structure: any $\mathbf{w}\in\mathcal{W}$ induces the event $\mathcal{C}$ with the same positive probability, whereas any assignment outside $\mathcal{W}$ induces $\mathcal{C}$ with probability zero. Therefore,
$p(\mathcal{C}\mid\mathbf{W}=\mathbf{w})\propto\mathbb{I}\{\mathbf{w}\in\mathcal{W}\},$
and hence
$p(\mathbf{W}\mid\mathcal{C})\propto p(\mathcal{C}\mid\mathbf{W})\,p(\mathbf{W})\propto\mathbb{I}\{\mathbf{W}\in\mathcal{W}\}\,p(\mathbf{W}),$
which is exactly the conditional randomization distribution used in (7).